# The interplay of emission geometry and extragalactic magnetic fields in the spectra of ultra-high-energy cosmic rays

### v1

@@ Abstract
The origins and acceleration mechanisms of ultra-high-energy cosmic rays (UHECRs) are unknown. Many models attribute their extreme energies to powerful astrophysical jets. Understanding whether jet geometry – specifically the opening angle and its orientation relative to Earth – affects observational signatures, is crucial for interpreting UHECR data. In this work, we perform numerical simulations of UHECR propagation in a magnetized universe to investigate the spectral signatures of jetted and non-jetted astrophysical sources. We demonstrate, for the first time, that under certain conditions, emission geometry can play a decisive role in shaping the observed spectrum of individual UHECR sources. These findings provide new insights into the conditions necessary for detecting UHECRs from jets, and highlight how the interplay between emission geometry and magnetic fields influences observed energy spectra.
@@ I Introduction
[…]
@@
The mechanisms behind the acceleration of UHECRs are still being investigated. Among the candidate sources of these particles are the astrophysical jets. These jets are often supersonic, so can produce strong shock waves. They create environments where velocity shear and turbulence are important, and also transport strong magnetic fields. Each of these effects can lead to the acceleration of high-energy particles. On this matter, many studies have been attempted in order to understand the precise mechanisms by which particles are accelerated within these jets. They suggest that these particles could be accelerated through various processes, each operating in a different region. Magnetic reconnection at the highly magnetized jet base [9], diffusive shock acceleration in regions such as the jet beam, backflow, and termination shocks [10], and shear acceleration in high-speed gradient regions [11] are possible contributors.
The acceleration processes and how cosmic rays interact with the magnetic fields in these environments remain complex. As UHECRs travel through the jet’s turbulent lobes, they may lose their directional information through scattering, but this depends on the scattering length and the coherence of the magnetic field. Although the interaction with turbulent magnetic fields and scattering in the lobes tends to isotropize the emission, certain acceleration mechanisms can cause the cosmic rays to retain a directional preference. For instance, UHECRs accelerated in relativistic blobs within the jet are expected to experience relativistic beaming, potentially resulting in anisotropic emission in the lab frame, depending on the observer’s viewpoint  [12]. This non-isotropic emission could offer crucial insights into both the acceleration mechanisms and the jet’s underlying geometry.
In this work, assuming that UHECRs are accelerated and emitted along the jet direction, we investigate the effect of the geometry of cosmic-ray emission and the extragalactic magnetic field properties along the propagation on the observed energy spectrum. In Section II, we briefly discuss CR diffusion in the context of a particle propagation in a turbulent magnetic field. Section III describes the simulations used in this work. The results are presented and discussed in Section IV. Finally, the conclusions are presented in Section VI.
@@ II Diffusive Cosmic Ray Propagation in Turbulent Magnetic Fields
Magnetic fields in galaxy clusters range from a few to tens of $\mu$G in their cores, down to $1-10$ nG in the outskirts [13, 14, 15, 16]. Estimates of magnetic fields in the filaments connecting galaxy clusters point to strengths of approximately $30-100$ nG [17, 18, 19]. Recent observational studies on magnetic fields in less dense regions, such as cosmic voids, estimate strengths ranging from $10^{-5}$ pG to $10^{-3}$ pG [20], although CMB constraints place an upper bound of $\sim 10\;\text{pG}$ [21]. The coherence length of void magnetic fields can vary enormously, with only weak constraints existing, mildly favoring values between 100 kpc and 100 Mpc [22]. In the Local Supercluster, estimates suggest that the root mean square (RMS) strength of magnetic fields lies between 1 nG and 100 nG [23, 24], with coherence lengths ranging from 10 kpc to 1 Mpc.
The Galactic magnetic field (GMF) is relatively strong compared to the extragalactic fields, reaching strengths of a few $\mu$[25, 26, 27, 28]. It can significantly deflect CRs arriving in the Galaxy. However, it does not significantly affect the propagation of UHECRs with energies above 1 EeV, since it acts on length scales much shorter than the typical interaction lengths. Therefore, it does not affect the CR spectrum measured on Earth.
Due to the limited knowledge and significant uncertainties regarding the structure and properties of extragalactic magnetic fields of large-scale environments, such as cosmic filaments, we simplify our study by focusing on the motion of CRs through a turbulent extragalactic magnetic field. This simplification assumes that the field can be fully characterized by its RMS strength, $B\equiv\sqrt{\left<B^{2}(x)\right>}$, and its coherence length, $L_{B}$, which defines the maximum spatial scale over which the magnetic field remains correlated. Although we acknowledge that this may noy be a fully accurate approximation, our approach captures the dominant features of turbulent scattering, being a reasonable framework for modeling CR propagation.
When a cosmic ray propagates through a turbulent extragalactic magnetic field, it may undergo multiple deflections. This behavior can be treated as a random walk, allowing the definition of a diffusion coefficient and an associated diffusion length, which corresponds to the average distance a particle travels before experiencing a significant deflection, providing a framework to describe its motion. The deviation of the particle’s trajectory after each deflection depends on its energy, charge, and the strength and coherence length of the magnetic field. To facilitate the analysis of diffusive cosmic ray propagation in such environments, a characteristic energy scale, known as the critical energy ($E_{c}$), is introduced.
The critical energy $E_{c}$ is defined as the energy for which the particle’s Larmor radius ($r_{L}$),
$r_{L}=\dfrac{E}{ZeB}\approx\dfrac{1.1}{Z}\left(\dfrac{E}{\text{EeV}}\right)\left(\dfrac{B}{\text{nG}}\right)^{-1}\;\text{Mpc}\,,$ (1)
@@
$E_{c}=ZeBL_{B}\approx 0.9Z\ \left(\dfrac{B}{\rm nG}\right)\ \left(\dfrac{L_{B}}{\rm Mpc}\right)\;{\rm EeV}\,.$ (2)
The diffusion coefficient $D(E)$ is given by $D(E)=cl_{D}/3$, where $l_{D}$ is the diffusion length. Ref. [29] proposed a fit for the diffusion coefficient that closely matches the numerical integration results:
$D(E)\simeq\frac{cL_{B}(z)}{3}\left[4\left(\frac{E}{E_{c}}\right)^{2}+a_{I}\left(\frac{E}{E_{c}}\right)+a_{L}\left(\frac{E}{E_{c}}\right)^{\alpha}\right].$ (3)
@@
$l_{D}(E)\simeq L_{B}\left[4\left(\frac{E}{E_{c}}\right)^{2}+a_{I}\left(\frac{E}{E_{c}}\right)+a_{L}\left(\frac{E}{E_{c}}\right)^{\alpha}\right],$ (4)
@@
Below the critical energy $E_{c}$, the Larmor radius is smaller than the coherence length and the particle undergoes stronger deflections. As a result, the flux from a given source becomes suppressed due to magnetic horizon effects [31]. This is quantified by the diffusion length $l_{D}$, in this regime $l_{D}\approx L_{B}(E_{c}/E)^{\alpha}$. For energies above $E_{c}$, the Larmor radius exceeds the coherence length, and the cosmic rays experience non-resonant scattering, yielding $l_{D}\approx L_{B}(E_{c}/E)^{2}$, indicating a rapid increase of diffusion length with energy. Therefore, if $l_{D}$ is much smaller than the distance between the source and the observer $R$, spatial diffusion occurs; if $l_{D}>R$, the particles propagate quasi-rectilinearly.
@@ III Simulations
where $\gamma$ is the spectral index of the source, $f_{\text{cut}}$ is the cut-off function, which depends on the charge ($Ze$) of the CR nucleus. We also assume that the cosmic-ray luminosity is constant over time.
The magnetic field is assumed to have a Kolmogorov power spectrum, with coherence length $L_{B}$ and an RMS field strength $B$. It is stored in a three-dimensional cubic grid of 100 kpc cell spacing, within $10$ Mpc box that repeats itself periodically. We perform the simulations considering several different realizations of the magnetic field, to reduce fluctuations and to properly capture the statistical properties of the problem.
Since the size of our galaxy is negligible compared to the vastness of the universe, performing a 3D forward simulation of particle propagation from a specific extragalactic location to Earth, is extremely time-consuming and nearly impossible to achieve. To overcome this issue and to estimate the observed energy spectrum from a given source located at a distance $R$, emitting particles in a jet with an opening angle of $\Psi$, we use the large sphere approach from Ref. [34]. First, cosmic rays are emitted isotropically from a single source. Particles are propagated and collected when they cross the surface of a sphere with radius $R$. This is done in CRPropa by defining a observer as a sphere of radius $R$ centered at the source location. Figure 1 illustrates the geometry considered in the simulation. For each simulated particle, we record its initial momentum $\vec{p}_{0}$, energy $E_{0}$, and particle identification $(A_{0},Z_{0})$, as well as its final energy $E_{f}$, particle identification $(A_{f},Z_{f})$ and detection position $\vec{r}_{det}$, which is the point where the particle intersects the surface of the detection sphere. The probability (and consequently the energy spectrum) of detecting a given CR emitted within a jet arriving at Earth with $E_{f}$ depends on the angle $\theta$ between the jet direction and Earth-source line of sight. We account for this considering that each detection position $\vec{r}_{det}$ emulates the Earth position. Therefore, for each detected event, we can compute the angle $\phi$ between the line of sight from the source to the Earth and the particle’s emission direction, which is given by
@@
Finally, events satisfying $\theta-\Psi/2\leq\phi\leq\theta+\Psi/2$ are considered to represent particles arriving at Earth from a jet with an opening angle $\Psi$ oriented at an angle $\theta$ relative to the Earth–source line of sight.
Figure 1: Illustration of the simulation setup. Particles are injected isotropically from the source (yellow star), and are collected as they cross the surface of a sphere with radius $R\equiv\left|\vec{r}_{\text{det}}\right|$. Each detection position $\vec{r}_{\text{det}}$ represents a possible Earth-like observer location. A simulated event with initial momentum $\vec{p}_{0}$ (dotted purple line) that satisfy the angular condition $\theta-\Psi/2\leq\phi\leq\theta+\Psi/2$ is considered to originate from a jet with opening angle $\Psi$, oriented at an angle $\theta$ relative to the line of sight between the source and the Earth. The orange line illustrates the trajectory of an event.
@@ IV Results
[…]
@@
  • magnetic field strength ($B$): $10^{-8}\;\text{G}$, $10^{-9}\;\text{G}$, $10^{-10}\;\text{G}$;
[…]
@@
In section IV.1 we present the results for our benchmark scenario, assuming proton primaries. This is followed by section IV.2, where we present results for nitrogen nuclei. In sections IV.3 and  IV.4 we discuss the dependence of the results on the spectral index and on the maximal rigidity of the sources, respectively. Section IV.5 directly addresses the emission geometry, comparing the isotropic and jetted cases.
IV.1 The pure-proton scenario
@@
First, we present the results of simulations involving $10^{5}$ protons. In figure 2, they are presented in terms of the modification factor $\xi$ (see Eq. 7). The columns correspond to different values of $R$: the first column represents $R=10\;\text{Mpc}$, the second column corresponds to $R=40\;\text{Mpc}$, and the third column to $R=100\;\text{Mpc}$. Similarly, the rows indicate different values of $\theta$: the first row represents $\theta=0^{\circ}$, the second corresponds to $\theta=40^{\circ}$, and the third to $\theta=80^{\circ}$. The different line styles represent cases where the magnetic field strength is $B=0.1\;\text{nG}$ (dashed-dotted line), $B=1.0\;\text{nG}$ (continuous line) and $B=10.0\;\text{nG}$ (dashed line). The purple color represents the results for a coherence length of $L_{B}=0.1\;\text{Mpc}$ while red curves show results for $L_{B}=1.0\;\text{Mpc}$, and yellow curves for $L_{B}=10.0\;\text{Mpc}$. The critical energy, defined by Eq. 2, is marked by a vertical line, with its color and style corresponding to the specified coherence length and magnetic field strength. In certain panels, they are not visible because their values fall outside the displayed range of the figure. In others, there was a superposition among the lines due to the combination of the parameter values. To fix that and improve clarity, an artificial small displacement of $0.1$ was introduced, the Ec values when $B=0.1\;\text{nG}$ were multiplied by an $0.9$ factor and the Ec values corresponding to $B=10.0\;\text{nG}$ by a $1.1$ factor.
@@
Figure 2: Modification factor obtained for proton primaries. Each column correspond to a different source–Earth distance ($R$), where $R=$ 10, 40, and 100 Mpc, respectively, from left to right. The rows correspond to different viewing angles ($\theta$), namely $\theta=0^{\circ}$, $40^{\circ}$, and $80^{\circ}$, from top to bottom. The jet opening angle is assumed to be $\Psi=15^{\circ}$. Line styles indicate different magnetic field strengths ($B$), while line colors correspond to different coherence lengths, $L_{B}$. Vertical lines mark the critical energy, defined as in eq. 2, with the color and line style matching the respective coherence length and magnetic field.
@@
The observed peaks in the figures, with $\xi$ reaching values greater than 1 for $\theta=0^{\circ}$ (first row), are attributed to pile-up effects caused by energy losses due to photopion interactions. Notably, the peak becomes more pronounced as the source distance increases, which reflects the higher probability of photon-pion interactions at greater distances. Concerning the magnetic field of $B=0.1\;\text{nG}$ (dashed-dotted line), it can also be observed that $\xi$ decreases at lower energies, as some particles undergo magnetic deflections larger than $\Psi/2=15^{\circ}$ and do not reach Earth when the jet direction is aligned with Earth. Since the deflection increases with the path length, the energy at which $\xi$ deviates from 1 shifts to higher values as the source–Earth distance ($R$) increases for the same parameter values ($B$ and $L_{B}$). Concerning the cases where the jet is misaligned with Earth (second and third rows) and the magnetic field strength is $B=0.1\;\text{nG}$, $\xi$ approaches zero as the energy increases. This behavior occurs because the magnetic field strength is insufficient to deflect these particles coming from jet angles of $\theta=40^{\circ}$ or $\theta=80^{\circ}$. However, in these cases, the curves shift to higher energies as the source distance increases, since the angular deflection becomes larger as particles describe longer trajectories. Since the critical energy is proportional to $L_{B}$, for a given energy and magnetic field strength $B$, particle deflection becomes more pronounced for larger values of $L_{B}$, as the diffusive regime begins to dominate over the ballistic. This explains why the right side of the curves for the misaligned jets for smaller coherence lengths approaches zero at energies lower than those with higher $L_{B}$ values. In addition, this is also the reason why the deflection at lower energies is more pronounced for the aligned jet ($\theta=0^{\circ}$) as $L_{B}$ increases.
@@
Considering a magnetic field strength of $B=1.0\;\text{nG}$ (continuous line), the energy at which $\xi$ falls below 1 shifts to higher values for $\theta=0^{\circ}$ (first row), since magnetic deflection is greater compared to the case with $B=0.1\;\text{nG}$ (dashe-dotted line). In other words, only the most energetic particles can remain deflected by less than $\Psi/2$ during propagation and still reach Earth. On the other hand, for $\theta=40^{\circ}$ and $\theta=80^{\circ}$ (second and third row), the endpoints of the corresponding curves also shift to higher energies, as more energetic particles are able to arrive at Earth, despite experiencing stronger magnetic deflection under $B=1.0\;\text{nG}$.
The same effects discussed for magnetic field magnitudes of $B=0.1\;\text{nG}$ and $B=1\;\text{nG}$ are also present and amplified in the $B=10\;\text{nG}$ case. For $\theta=0^{\circ}$, the energy at which $\xi$ falls below 1 is shifted to higher energies, while the right side of the curves corresponding to $\theta=40^{\circ}$ and $\theta=80^{\circ}$ are also shifted to the right. However, for a source distance of $R=100\;\text{Mpc}$, the number of protons in the lower energy range decreases and $\xi$ approaches zero as a consequence of the magnetic horizon. This is explained by the long propagation distance of these events in such a highly diffusive regime, resulting in a propagation time longer than the age of the Universe.
This can be confirmed by analyzing Fig. 3, where the distribution of events is plotted as a function of energy and trajectory elongation ($d$). The trajectory elongation is defined here as the difference between the distance traveled by the particle and the source–Earth distance:

### v2

# Effects of extragalactic magnetic field on the spectra of ultra-high-energy cosmic rays from jetted sources

@@ Abstract
The origins and acceleration mechanisms of ultra-high-energy cosmic rays (UHECRs) are unknown. Many models attribute their extreme energies to powerful astrophysical jets. Understanding whether jet geometry – specifically the opening angle and its orientation relative to Earth – affects observational signatures is crucial for interpreting UHECR data. In this work, we perform numerical simulations of UHECR propagation in a magnetized universe to investigate the spectral signatures of jetted and nonjetted astrophysical sources. We demonstrate, for the first time, that under certain conditions, emission geometry can play a decisive role in shaping the observed spectrum of individual UHECR sources. These findings provide new insights into the conditions necessary for detecting UHECRs from jets, and highlight how the interplay between emission geometry and magnetic fields influences observed energy spectra.
@@ I Introduction
UHECR are particles with energies above $\text{EeV}11$11 EeV $\equiv$ $10^{18}\;\text{eV}$. that hit our atmosphere coming from space. Since their discovery in 1963 [1], their origins and the mechanisms capable of accelerating them to such extraordinary energies remain unknown, posing one of the most challenging mysteries in modern astrophysics [2, 3].
@@
The mechanisms behind the acceleration of UHECRs are still being investigated. Among the candidate sources of these particles are the astrophysical jets. These jets are often supersonic, so they can produce strong shock waves. They create environments where velocity shear and turbulence are important and also transport strong magnetic fields. Each of these effects can lead to the acceleration of high-energy particles. On this matter, many studies have been attempted in order to understand the precise mechanisms by which particles are accelerated within these jets. They suggest that these particles could be accelerated through various processes, each operating in a different region. Magnetic reconnection at the highly magnetized jet base [9], diffusive shock acceleration in regions such as the jet beam, backflow, and termination shocks [10], and shear acceleration in high-speed gradient regions [11] are possible contributors.
The acceleration processes and how cosmic rays interact with the magnetic fields in these environments remain complex. As UHECRs travel through the jet’s turbulent lobes, they may lose their directional information through scattering, but this depends on the scattering length and the coherence of the magnetic field. Although the interaction with turbulent magnetic fields and scattering in the lobes tends to isotropize the emission, certain acceleration mechanisms can cause the cosmic rays to retain a directional preference. For instance, UHECRs accelerated in relativistic blobs within the jet are expected to experience relativistic beaming, potentially resulting in anisotropic emission in the laboratory frame, depending on the observer’s viewpoint  [12]. This nonisotropic emission could offer crucial insights into both the acceleration mechanisms and the jet’s underlying geometry.
In this work, assuming that UHECRs are accelerated and emitted along the jet direction, we investigate the effect of the geometry of cosmic-ray emission and the extragalactic magnetic field properties along the propagation on the observed energy spectrum. In Sec. II, we briefly discuss CR diffusion in the context of a particle propagation in a turbulent magnetic field. Section III describes the simulations used in this work. The results are presented and discussed in Sec. IV. Finally, the conclusions are presented in Sec. VI.
@@ II Diffusive Cosmic Ray Propagation in Turbulent Magnetic Fields
[…]
The Galactic magnetic field (GMF) is relatively strong compared to the extragalactic fields, reaching strengths of a few $\mu$[26, 27, 28, 29]. However, it does not significantly affect the CR spectrum measured on Earth.
Due to the limited knowledge and significant uncertainties regarding the structure and properties of extragalactic magnetic fields of large-scale environments, such as cosmic filaments, we simplify our study by focusing on the motion of CRs through a turbulent extragalactic magnetic field. This simplification assumes that the field can be fully characterized by its RMS strength, $B\equiv\sqrt{\left<B^{2}(x)\right>}$, and its coherence length, $L_{B}$, which defines the maximum spatial scale over which the magnetic field remains correlated. Although we acknowledge that this may not be a fully accurate approximation, it still captures the dominant features of turbulent scattering and thus provides a reasonable framework for modeling CR propagation.
When a cosmic ray propagates through a turbulent EGMF, it may undergo multiple deflections. This behavior can be treated as a random walk, allowing the definition of a diffusion coefficient and an associated diffusion length, which corresponds to the average distance a particle travels before experiencing a significant deflection, providing a framework to describe its motion. The deviation of the particle’s trajectory after each deflection depends on its energy, charge, and the strength and coherence length of the magnetic field. To facilitate the analysis of diffusive cosmic-ray propagation in such environments, a characteristic energy scale, known as the critical energy ($E_{\text{c}}$), is introduced.
The critical energy $E_{\text{c}}$ is defined as the energy for which the particle’s Larmor radius ($r_{\text{L}}$),
$r_{\text{L}}=\dfrac{E}{ZeB}\approx\dfrac{1.1}{Z}\left(\dfrac{E}{\text{EeV}}\right)\left(\dfrac{B}{\text{nG}}\right)^{-1}\;\text{Mpc}\,$, (1)
@@
$E_{\text{c}}=ZeBL_{B}\approx 0.9Z\ \left(\dfrac{B}{\rm nG}\right)\ \left(\dfrac{L_{B}}{\rm Mpc}\right)\;{\rm EeV}\,.$ (2)
The diffusion coefficient $D(E)$ is given by $D(E)=cl_{\text{D}}/3$, where $l_{\text{D}}$ is the diffusion length. Reference [30] proposed a fit for the diffusion coefficient that closely matches the numerical integration results:
$D(E)\simeq\frac{cL_{B}(z)}{3}\left[4\left(\frac{E}{E_{\text{c}}}\right)^{2}+a_{I}\left(\frac{E}{E_{\text{c}}}\right)+a_{L}\left(\frac{E}{E_{\text{c}}}\right)^{\alpha}\right].$ (3)
@@
$l_{\text{D}}(E)\simeq L_{B}\left[4\left(\frac{E}{E_{\text{c}}}\right)^{2}+a_{I}\left(\frac{E}{E_{\text{c}}}\right)+a_{L}\left(\frac{E}{E_{\text{c}}}\right)^{\alpha}\right],$ (4)
@@
Below the critical energy $E_{\text{c}}$, the Larmor radius is smaller than the coherence length and the particle undergoes stronger deflections. As a result, the flux from a given source becomes suppressed due to magnetic horizon effects [32]. This is quantified by the diffusion length $l_{\text{D}}$, in this regime $l_{\text{D}}\approx L_{B}(E_{\text{c}}/E)^{\alpha}$. For energies above $E_{\text{c}}$, the Larmor radius exceeds the coherence length, and the cosmic rays experience nonresonant scattering, yielding $l_{\text{D}}\approx L_{B}(E_{\text{c}}/E)^{2}$, indicating a rapid increase of diffusion length with energy. Therefore, if $l_{\text{D}}$ is much smaller than the distance between the source and the observer $R$, spatial diffusion occurs; if $l_{\text{D}}>R$, the particles propagate quasirectilinearly.
@@ III Simulations
where $\gamma$ is the spectral index of the source and $f_{\text{cut}}$ is the cutoff function, which depends on the charge ($Ze$) of the CR nucleus. We also assume that the cosmic-ray luminosity is constant over time.
The magnetic field is assumed to have a Kolmogorov power spectrum, with coherence length $L_{B}$ and a RMS field strength $B$. It is stored in a three-dimensional cubic grid of 10 kpc cell spacing, within $10$ Mpc box that repeats itself periodically. We perform the simulations considering several different realizations of the magnetic field, to reduce fluctuations and to properly capture the statistical properties of the problem.
Since the size of our Galaxy is negligible compared to the vastness of the Universe, performing a three-dimensional forward simulation of particle propagation from a specific extragalactic location to Earth is extremely time consuming and nearly impossible to achieve. To overcome this issue and to estimate the observed energy spectrum from a given source located at a distance $R$, emitting particles in a jet with an opening angle of $\Psi$, we use the large sphere approach from Ref. [35]. First, cosmic rays are emitted isotropically from a single source. Particles are propagated and collected when they cross the surface of a sphere with radius $R$. This is done in CRPropa by defining an observer as a sphere of radius $R$ centered at the source location. Figure 1 illustrates the geometry considered in the simulation. For each simulated particle, we record its initial momentum $\vec{p}_{0}$, energy $E_{0}$, and particle identification $(A_{0},Z_{0})$, as well as its final energy $E_{f}$, particle identification $(A_{f},Z_{f})$, and detection position $\vec{r}_{det}$, which is the point where the particle intersects the surface of the detection sphere. The probability (and consequently the energy spectrum) of detecting a given CR emitted within a jet arriving at Earth with $E_{f}$ depends on the angle $\theta$ between the jet direction and Earth-source line of sight. We account for this considering that each detection position $\vec{r}_{det}$ emulates the Earth position. Therefore, for each detected event, we can compute the angle $\phi$ between the line of sight from the source to the Earth and the particle’s emission direction, which is given by
@@
Finally, events satisfying $\theta-\Psi/2\leq\phi\leq\theta+\Psi/2$ are considered to represent particles arriving at Earth from a jet with an opening angle $\Psi$ oriented at an angle $\theta$ relative to the Earth-to-source line of sight [^2] Notice that the observed energy spectrum for $\theta\neq 0^{\circ}$ must be corrected by a geometric factor, given by the ratio between the cone solid angle and the ring solid angle..
Figure 1: Illustration of the simulation setup. Particles are injected isotropically from the source (yellow star), and are collected as they cross the surface of a sphere with radius $R\equiv\left|\vec{r}_{\text{det}}\right|$. Each detection position $\vec{r}_{\text{det}}$ represents a possible Earth-like observer location. A simulated event with initial momentum $\vec{p}_{0}$ (dotted purple line) that satisfies the angular condition $\theta-\Psi/2\leq\phi\leq\theta+\Psi/2$ is considered to originate from a jet with opening angle $\Psi$, oriented at an angle $\theta$ relative to the line of sight between the source and the Earth. The orange line illustrates the trajectory of an event.
@@ IV Results
In this section we present the simulation results. To structure the discussion, we fix a few reference parameters, which we later vary to properly discuss their impact. The reference spectral index is $\gamma=1$, and the injection spectrum given by Eq. 5 is assumed to be a Heaviside function of the (initial) energy, being 1 below $1\;\text{ZeV}$ and 0 above.
@@
  • magnetic field strength ($B$): $10^{-8}\;\text{G}$, $^{-9}\;\text{G}, 1010$$^{-10}\;\text{G}$;
  • coherence length ($L_{B}$): $100\;\text{kpc}$, $1\;\text{Mpc}$, $10\;\text{Mpc}$;
@@
In Sec. IV.1 we present the results for our benchmark scenario, assuming proton primaries. This is followed by Sec. IV.2, where we present results for nitrogen nuclei. In Sec. IV.3 and  IV.4, we discuss the dependence of the results on the spectral index and on the maximal magnetic rigidity [^3] The magnetic rigidity $R$ of a charged particle is defined as $R=E/Ze$, where $E$ is the particle’s energy, $Z$ is the charge number, and $e$ is the elementary charge. of the sources, respectively. Section IV.5 directly addresses the emission geometry, comparing the isotropic and jetted cases.
IV.1 Pure-proton scenario
@@
First, we present the results of simulations involving $10^{5}$ protons. In Fig. 2, they are presented in terms of the modification factor $\xi$ (see Eq. 7). The columns correspond to different values of $R$: the first column represents $R=10\;\text{Mpc}$, the second column corresponds to $R=40\;\text{Mpc}$, and the third column corresponds to $R=100\;\text{Mpc}$. Similarly, the rows indicate different values of $\theta$: the first row represents $\theta=0^{\circ}$, the second corresponds to $\theta=40^{\circ}$, and the third corresponds to $\theta=80^{\circ}$. The different line styles represent cases where the magnetic field strength is $B=0.1\;\text{nG}$ (dot-dashed line), $B=1.0\;\text{nG}$ (continuous line) and $B=10.0\;\text{nG}$ (dashed line). The magenta color represents the results for a coherence length of $L_{B}=0.1\;\text{Mpc}$, while green curves show results for $L_{B}=1.0\;\text{Mpc}$, and blue curves show results for $L_{B}=10.0\;\text{Mpc}$. The critical energy, defined by Eq. 2, is marked by a vertical line, with its color and style corresponding to the specified coherence length and magnetic field strength. In certain panels, they are not visible because their values fall outside the displayed range of the figure. In others, there was a superposition among the lines due to the combination of the parameter values. To fix that and improve clarity, an artificial small displacement of $0.1$ was introduced, the $E_{\text{c}}$ values when $B=0.1\;\text{nG}$ were multiplied by an $0.9$ factor and the $E_{\text{c}}$ values corresponding to $B=10.0\;\text{nG}$ by a $1.1$ factor.
The observed peaks in the figures, with $\xi$ reaching values greater than 1 for $\theta=0^{\circ}$ (first row), are attributed to pile-up effects caused by energy losses due to photopion interactions. Notably, the peak becomes more pronounced as the source distance increases, which reflects the higher probability of photon-pion interactions at greater distances.
@@
Figure 2: Modification factor obtained for proton primaries. Each column corresponds to a different source-Earth distance ($R$), where $R=$ 10, 40, and 100 Mpc, respectively, from left to right. The rows correspond to different viewing angles ($\theta$), namely $\theta=0^{\circ}$, $40^{\circ}$, and $80^{\circ}$, from top to bottom. The jet opening angle is assumed to be $\Psi=15^{\circ}$. Line styles indicate different magnetic field strengths ($B$), while line colors correspond to different coherence lengths, $L_{B}$. Vertical lines mark the critical energy, defined as in Eq. 2, with the color and line style matching the respective coherence length and magnetic field.
@@
Concerning the magnetic field of $B=0.1\;\text{nG}$ (dot-dashed line), it can also be observed that $\xi$ decreases at lower energies, as some particles undergo magnetic deflections larger than $\Psi/2=7.5^{\circ}$ and do not reach Earth when the jet direction is aligned with Earth. Since the deflection increases with the path length, the energy at which $\xi$ deviates from 1 shifts to higher values as the source-Earth distance ($R$) increases for the same parameter values ($B$ and $L_{B}$). Concerning the cases where the jet is misaligned with Earth (second and third rows) and the magnetic field strength is $B=0.1\;\text{nG}$, $\xi$ approaches zero as the energy increases. This behavior occurs because the magnetic field strength is insufficient to deflect these particles coming from jet angles of $\theta=40^{\circ}$ or $\theta=80^{\circ}$. However, in these cases, the curves shift to higher energies as the source distance increases, since the angular deflection becomes larger as particles describe longer trajectories. Since the critical energy is proportional to $L_{B}$, for a given energy and magnetic field strength $B$, particle deflection becomes more pronounced for larger values of $L_{B}$, as the diffusive regime begins to dominate over the ballistic. This explains why the right side of the curves for the misaligned jets for smaller coherence lengths approaches zero at energies lower than those with higher $L_{B}$ values. In addition, this is also the reason why the deflection at lower energies is more pronounced for the aligned jet ($\theta=0^{\circ}$) as $L_{B}$ increases.
@@
Considering a magnetic field strength of $B=1.0\;\text{nG}$ (continuous line), the energy at which $\xi$ falls below 1 shifts to higher values for $\theta=0^{\circ}$ (first row), since magnetic deflection is greater compared to the case with $B=0.1\;\text{nG}$ (dot-dashed line). In other words, only the most energetic particles can remain deflected by less than $\Psi/2$ during propagation and still reach Earth. On the other hand, for $\theta=40^{\circ}$ and $\theta=80^{\circ}$ (second and third row), the endpoints of the corresponding curves also shift to higher energies, as more energetic particles are able to arrive at Earth, despite experiencing stronger magnetic deflection under $B=1.0\;\text{nG}$.
The same effects discussed for magnetic field magnitudes of $B=0.1\;\text{nG}$ and $B=1\;\text{nG}$ are also present and amplified in the $B=10\;\text{nG}$ case. For $\theta=0^{\circ}$, the energy at which $\xi$ falls below 1 is shifted to higher energies, while the right side of the curves corresponding to $\theta=40^{\circ}$ and $\theta=80^{\circ}$ are also shifted to the right. At low energies, the distribution of particle arrival directions becomes isotropic due to magnetic diffusion, and in this limit the modification factors converge to the ratio between the solid angle of a cone with aperture $\psi$ and that of a full sphere $\sim 4\times 10^{-3}$. However, for a source distance of $R=100\;\text{Mpc}$, the number of protons in the lower energy range decreases and $\xi$ approaches zero as a consequence of the magnetic horizon. This is explained by the long propagation distance of these events in such a highly diffusive regime, resulting in a propagation time longer than the age of the Universe.
This can be confirmed by analyzing Fig. 3, where the distribution of events is plotted as a function of energy and trajectory elongation ($d$). The trajectory elongation is defined here as the difference between the distance traveled by the particle and the source-Earth distance: